(D^3+1) y= 3+e^(-x) +5e^2x?

1
Expert's answer
2022-04-04T16:32:29-0400

"y'''+y=3+e^{-x}+5e^{2x}"

The solutions to a nonhomogeneous equation are of the form

"y(x) = y_c(x) + y_p(x)" ,

where "y_c" is the general solution to the associated homogeneous equation and "y_p" is a particular solution.

The associated homogeneous equation:

"y'''+y=0"

The general solution of this equation is determined by the roots of the characteristic equation:

"\lambda^3+1=0"

"(\lambda+1)(\lambda^2-\lambda+1)=0"

"\lambda_1=-1"

"(\lambda^2-\lambda+1)=0"

"D=b^2-4ac=1-4=-3"

"\lambda_{2,3}=\frac{1\pm i\sqrt 3}{2}" , where "i=\sqrt{-1}" .

"y_c(x)=c_1e^{-x}+e^{\frac x2}(c_2\cos{\frac{\sqrt3}{2}x}+c_3\sin{\frac{\sqrt3}{2}x})" .

The particular solution of the differential equation:

"y_p(x)=A+xBe^{-x}+Ce^{2x}"

"y_p'=Be^{-x}-xBe^{-x}+2Ce^{2x}"

"y_p''=-2Be^{-x}+xBe^{-x}+4Ce^{2x}"

"y_p'''=3Be^{-x}-xBe^{-x}+8Ce^{2x}"

Now put these into the original differential equation to get:

"3Be^{-x}-xBe^{-x}+8Ce^{2x}+A+xBe^{-x}+Ce^{2x}=""3+e^{-x}+5e^{2x}"

Equating coefficients, we get

"A=3";   "B=\frac13"; "C=\frac59".

"y_p(x)=3+\frac x3e^{-x}+\frac59e^{2x}"

Answer: "y=c_1e^{-x}+e^{\frac x2}(c_2\cos{\frac{\sqrt3}{2}x}+c_3\sin{\frac{\sqrt3}{2}x})+""3+\frac x3e^{-x}+\frac59e^{2x}" .